Walk-up ramp moves horizontally 20 ft while rising 5 ft. What is it's approximate length to the nearest tenth of a foot.

Answer by Boreal(15235)

Pythagorean theorem=5^2+20^2=length of ramp squared
L=sqrt(425 ft^2)=20.6 feet